OpenStudy (anonymous):

Write an equation in point-slope form of the line having the given slope that contains the given point M=5 (3,-2)

OpenStudy (anonymous):

ohhhhhh that makes it so much easier thanks!!!!

OpenStudy (alexandervonhumboldt2):

he is wrong

OpenStudy (alexandervonhumboldt2):

point slope form is y-y1=m*(x-x1)

OpenStudy (alexandervonhumboldt2):

m=5

OpenStudy (alexandervonhumboldt2):

then the equation is y+2=5*(x-3)

OpenStudy (anonymous):

are you sure

OpenStudy (alexandervonhumboldt2):

yeah

OpenStudy (trojanpoem):

Both of them are right dude, you can use either amorfide way or AlexandervonHumboldt2 method

OpenStudy (alexandervonhumboldt2):

yeah but they ask for point-slpe form

OpenStudy (anonymous):

ok so who is right

OpenStudy (amorfide):

alex

OpenStudy (anonymous):

ok thanks